# Generating Fock-state superposition from coherent state by quantum measurement

### v1

@@ III.1 Fock-state generation
Our protocol to generate an arbitrary Fock state $\left|n\right\rangle$ can be performed with a sequence of projective measurements $M_{e}=|e\rangle\langle e|$ on the excited state of ancillary qubit. In the beginning of the each round of evolution and measurement, the qubit starts from $\left|e\right\rangle$. The initial density matrix of the full system reads
$\rho(0)=|e\rangle\langle e|\otimes|\alpha\rangle\langle\alpha|,\quad\left|\alpha\right\rangle=\sum_{k=0}^{\infty}\alpha_{k}\left|k\right\rangle,$ (12)
where $\alpha_{k}=e^{-|\alpha|^{2}/2}\alpha^{k}/\sqrt{k!}$. On substituting Eq. (5) into Eq. (9), the diagonal part of density matrix of the resonator after $N$ rounds of equally-spaced evolution and measurement turns out to be
$\mathcal{D}[\rho_{a}(N\tau)]=\sum_{k=0}^{\infty}\dfrac{|\alpha_{k}|^{2}\cos^{2N}\Omega^{(e)}_{k}\tau|k\rangle\langle k|}{P_{e}(N)}.$ (13)
The fidelity of the target Fock state $\left|n\right\rangle$ is found to be
$\mathcal{F}(N)=\left\langle n\right|\rho_{a}(N\tau)\left|n\right\rangle=\frac{|\alpha_{n}|^{2}\cos^{2N}\Omega_{n}^{(e)}\tau}{\sum_{k}|\alpha_{k}|^{2}\cos^{2N}\Omega^{(e)}_{k}\tau}.$ (14)
To generate the Fock state $\left|n\right\rangle$, the relevant population-reduction factor has to satisfy $|\lambda^{(e)}_{n}(l\tau_{n}^{(e)})|^{2}=1$. Due to Eq. (5) and the resonant assumption $\delta=0$, the evolution period of a single round $\tau$ can be found as
$\tau=l\tau_{n}^{(e)}=l\frac{\pi}{\Omega^{(e)}_{n}}=l\frac{\pi}{g\sqrt{n+1}},$ (15)
where $l\subseteq\mathbb{N}+$. If the irrelevant population-reduction factors $|\lambda^{(e)}_{k}(l\tau_{k}^{(e)})|^{2}$ for all the significantly occupied Fock states $|k\rangle$, $k\neq n$, are always less than unit, the resonator state will gradually approach the desired $|n\rangle$ as the round number $N$ increases. The filter effect can be further amplified by the success probability $P_{e}(N)<1$ in the denominator of Eq. (13). Therefore the density matrix of the resonator mode becomes
@@
in the ideal large-$N$ limit. Also it indicates that the success probability converges to the initial population of the target Fock-state:
@@
Besides the target state $|n\rangle$, also the populations on the Fock states $|n_{j}^{(l)}\rangle$ remain invariant when
$n_{j}^{(l)}=\frac{j^{2}}{l^{2}}(n+1)-1$ (18)
with $j\subseteq\mathbb{N}+$ due to Eqs. (5) and (15). The solution set with $l\neq 1$ covers that with $l=1$. To perform strategies for Fock-state generation with varying $\tau$ or $l$ in the sequence of measurements, we then focus on the intersection of the solution set with $l=1$ and that with $l\neq 1$. Thus in the following, $n_{j}^{(l)}$ can be briefly labelled as $n_{j}$ and it is readily to find that $n=n_{1}$. More generally, a sufficiently large number of evolution-measurement cycles with $\tau$ in Eq. (15) give rise to a stabilized subspace $\mathcal{V}_{n}^{(e)}$ spanned by $\{\left|n_{j}\right\rangle|j\subseteq\mathbb{N}+\}$, where both the populations and coherence elements are under protection in our protocol. To enhance the final success probability of the target Fock state, the initial coherent state $\left|\alpha\right\rangle$ of the resonator can be prepared with $|\alpha|^{2}=n$ such that $|n\rangle$ is the maximally occupied Fock state and the populations on $|n_{j\geq 2}\rangle$ are negligible. For example, when $n=n_{1}=5$, the population on the next unwanted yet protected Fock-state $n_{2}=23$ is about $p_{n_{2}}\sim 10^{-5}$.
@@
Figure 2: Population-reduction factors as a function of the Fock index $k$ after $N$ evolution-measurement cycles with various evolution periods $\tau$ indicated by $l$. The target states are (a) $|n=5\rangle$ and (b) $|n=10\rangle$. The black solid lines indicate a single round with $\tau=\tau_{n}^{(e)}$ ($l=1$). The blue dotted lines, the red dashed lines, and the yellow dash-dotted lines indicate $N=5$ rounds with $\tau=\tau_{n}^{(e)}$, $\tau=2\tau_{n}^{(e)}$ ($l=2$), and $\tau=3\tau_{n}^{(e)}$ ($l=3$), respectively. $g=0.05\omega_{e}$ and $\delta=0$.
@@
The efficiency for generating the target state is determined by the profile of the population-reduction factors $|\lambda^{(e)}_{k}(l\tau_{n})|^{2N}$ in the Fock space as shown in Figs. 2(a) and 2(b), where $n=5$ and $n=10$, respectively. These factors are evaluated under various periods $\tau$ of free-evolution (indicated by $l$) after $N$ rounds of evolution and measurement. The black solid lines in Fig. 2 demonstrate the effect under a single projective measurement, confirming that the population-reduction factors of the irrelevant Fock states $k\neq n_{j\geq 1}$ satisfy $|\lambda_{k}^{(e)}(\tau)|^{2}<1$. And the comparison between the blue dotted lines and the black solid lines suggests that the populations in the neighboring regions of $|n_{j}\rangle$ can be significantly suppressed by only $5$ measurements.
@@
Using multiples of $\tau$ (see the red dashed lines for $l=2$ and the yellow dot-dashed line for $l=3$), it is interesting to find that on one hand, they can intensify the population-reduction effect surrounding the target state $|n_{1}\rangle$ with a much concentrated population-protected region; yet on the other hand, more population-protected (although not strictly) Fock states arise, e.g., see the high profiles between $|n_{1}=5\rangle$ and $|n_{2}=23\rangle$ in Fig. 2(a). Those results suggest a compromise or hybrid strategy labelled with $\mathcal{S}_{l}^{(q)}$ that filters out the populations distributed on the states other than $|n_{j}\rangle$ by using $\tau=\tau^{(e)}_{n}$ in the first $q$ rounds of evolution-and-measurement and then promotes the state-generation efficiency by using $\tau=l\tau^{(e)}_{n}$, $l>1$, in the rest rounds. Often we take $l=2$ or $l=3$ since the advantage from an even larger $l$ is not significant. Accordingly, the uniform strategy with an invariant period $\tau=\tau^{(e)}_{n}$ can be labelled with $\mathcal{S}_{1}^{(\infty)}$.
@@
Figure 3: Comparison of various strategies in Fock-state generation of (a) and (c): $|5\rangle$ and of (b) and (d): $|10\rangle$ as a function of the measurement number $N$. (a) and (b): Target-state fidelity. (c) and (d): Success probability, where the gray dotted lines indicate the initial population of $|n\rangle$. The system parameters are the same as Fig. 2.
@@
In Fig. 3, we compare the efficiencies for generating the Fock state between the uniform strategy and the hybrid strategy $\mathcal{S}^{(5)}_{l}$ with $l=2,3$ in terms of the target-state fidelity $\mathcal{F}(N)$ and the success probability $P_{e}(N)$, whose definitions can be found in Eqs. (14) and (11), respectively. With the current hybrid strategy, the first $q=5$ rounds of the evolution and measurement use $\tau=\tau_{n}^{(e)}$ and the rest rounds use $\tau=l\tau_{n}^{(e)}$. It is found that the hybrid strategy overwhelms the uniform one in the state fidelity under a limited number of measurements. And a larger $l$ gives rise to an even better performance. In Fig. 3(a), the fidelity $\mathcal{F}$ of the target Fock-state $\left|5\right\rangle$ can reach $99.6\%$ within $15$ measurements when $l=3$. In contrast, $\mathcal{F}$ reaches only $68.6\%$ within $20$ measurements by the uniform strategy. In Fig. 3(b), it is found that a higher Fock state $\left|10\right\rangle$ could be obtained by merely $30$ measurements in the strategy $\mathcal{S}^{(5)}_{3}$ with a fidelity over $98.4\%$. Different strategy yields the same asymptotic value of the success probability, i.e., the population of $|n\rangle$ in the initial coherent state, which can be obtained by Eq. (17) with $|\alpha|^{2}=n$. In both Figs. 3(c) and 3(d), a larger $l$ gives rise to a faster convergence.
@@
Figure 4: Dependence of (a) Fidelity and (b) Success probability on the measurement number $N$ under the strategy $\mathcal{S}_{3}^{(5)}$ for various target state. The system parameters are the same as Fig. 2.
@@
In Fig. 4, we compare the performance for generating different Fock states under the hybrid strategy $\mathcal{S}_{3}^{(5)}$. And the initial coherent states are accordingly set with $|\alpha|^{2}=n$. It is found in Fig. 4(a) that more rounds of evolution-measurement cycles are required to prepare a higher Fock states. That could be understood that a larger $n$ indicates a broader population distribution in Fock space and a smaller overlap between the corresponding coherent state and the target state. For example, it requires only $6$ rounds of evolution and measurement to enhance the fidelity of the target Fock state $\left|2\right\rangle$ to $99\%$, yet about $23$ rounds to prepare $\left|8\right\rangle$ with almost the same fidelity. Figure 4(b) shows that the behavior of the success probability is insensitive to $n$. It declines rapidly for the first few rounds and approaches the steady value after about $N=10$ rounds.
@@ III.2 Superposed Fock-state generation
Beyond the existing works for Fock state generation, our protocol can be used to prepare a superposed Fock state, such as
@@
with desired $n$, $c_{0}$, and $c_{n}$. This target can be achieved by preparing the ancillary qubit in its ground state $|g\rangle$ and performing a sequence of projective measurements $M_{g}=|g\rangle\langle g|$ on the ancillary qubit. Due to Eqs. (4) and (5), the effective time-evolution operator for the resonator under the resonant condition $\delta=0$ and an invariant $\tau$ is now written as,
@@
In this case, the population on the vacuum state $\left|0\right\rangle$ of the resonator is always under protection due to the decoupling of the state $\left|g0\right\rangle$ from the other composite states in the time evolution. Using Eqs. (5) and (9), the diagonal part of the density matrix of the resonator after $N$ rounds of equally-spaced evolution-and-measurement can be expressed by
@@
From Eqs. (19), (21), and (23), the fidelity $\mathcal{F}_{\pm}$ about the target state $\left|\psi_{\pm}\right\rangle$ under $N$ equally-spaced evolution-and-measurement rounds can be obtained as
$\mathcal{F}_{\pm}=$ $\bigg{[}|c_{0}|^{2}p_{0}+|c_{n}|^{2}p_{n}\cos^{2N}\left(\sqrt{k}g\tau\right)$
±$\pm 2{\rm Re}[C_{n0}c_{0}c_{n}^{*}]\cos^{N}\left(\sqrt{k}g\tau\right)\bigg{]}/P_{g}(N).$ (24)
Figure 5: (a) Fidelities $\mathcal{F}_{\pm}$ for the target superposed state $(\left|0\right\rangle\pm\left|5\right\rangle)/\sqrt{2}$ or their combination $\mathcal{F}_{+}+\mathcal{F}_{-}$ as functions of $N$ under the uniform strategy $\mathcal{S}_{1}^{(\infty)}$ with $\tau=\tau_{5}^{(g)}$ and a hybrid strategy $\mathcal{S}_{3}^{(5)}$ that modifies the round period to be $\tau=3\tau_{5}^{(g)}$ after $5$ rounds of evolution and measurement. (b) Success probability of finding the qubit in its ground state after $N$ measurements. The system parameters are the same as Fig. 2.
Both populations on the ground state $\left|0\right\rangle$ and the Fock state $\left|n\right\rangle$ and their coherence should be under protection to generate the target superposed state $|\psi_{\pm}\rangle$ defined in Eq. (19). To this end, the evolution periods for each round of evolution and measurement are selected as
@@
where $l\in\mathbb{N}^{+}$. With $l=1$, a stabilized subspace under a full protection emerges under multiple measurements. It reads,
$\mathcal{V}^{(g)}_{n}={\rm span}\{\left|0\right\rangle,\left|n\right\rangle,\left|2^{2}n\right\rangle,\left|3^{2}n\right\rangle,\cdots\}.$ (26)
The ratio of the probability amplitudes $c_{0}/c_{n}$ is determined by $\alpha_{0}/\alpha_{n}$. Then inversely, one can find that the initial coherent state of the resonator is $\left|\alpha\right\rangle$ with $\alpha=(c_{n}\sqrt{n!}/c_{0})^{1/n}$. Similar to the Fock-state generation described in Sec. III.1, the initial populations on the Fock-state elements in the stabilized states $\{\left|j^{2}n\right\rangle|j\geqslant 2\}$ are negligible. Consequently, the diagonal and off-diagonal parts of the density matrix of the resonator in the large $N$ limit read
$\mathcal{D}[\rho_{a}(N\tau)]$ $\rightarrow e^{-|\alpha|^{2}}\left(|0\rangle\langle 0|+\frac{|c_{n}|^{2}}{|c_{0}|^{2}}|n\rangle\langle n|\right)/P_{g}(N),$ (27)
$\mathcal{C}[\rho_{a}(N\tau)]$ $\rightarrow e^{-|\alpha|^{2}}\left[\frac{c_{n}}{c_{0}}(-1)^{N}|n\rangle\langle 0|+{\rm H.c.}\right]/P_{g}(N),$ (28)
@@
$P_{g}(N)\rightarrow p_{0}+p_{n}=\exp\left[-\left(\frac{|c_{n}|}{|c_{0}|}\sqrt{n!}\right)^{\frac{2}{n}}\right]/|c_{0}|^{2}.$ (29)
@@
Figure 6: Density matrix of the resonator after various rounds of evolution and measurement under (a) and (c) the uniform strategy with $\tau=\tau_{5}^{(g)}$ or (b) and (d) the hybrid strategy $\mathcal{S}_{3}^{(5)}$. In (a) and (b): $N=10$; in (c) and (d): $N=11$. The target state is $(\left|0\right\rangle\pm\left|5\right\rangle)/\sqrt{2}$ and other parameters are the same as in Fig. 2.

### v2

@@ III.1 Fock-state generation
Our protocol to generate an arbitrary Fock state $|n\rangle$ can be performed with a sequence of projective measurements $M_{e}=|e\rangle\langle e|$ on the excited state of ancillary qubit. In the beginning of the each round of evolution and measurement, the qubit starts from $|e\rangle$. The initial density matrix of the full system reads
$\rho(0)=|e\rangle\langle e|\otimes|\alpha\rangle\langle\alpha|,\quad|\alpha\rangle=\sum_{k=0}^{\infty}\alpha_{k}|k\rangle,$ (12)
where $\alpha_{k}=e^{-|\alpha|^{2}/2}\alpha^{k}/\sqrt{k!}$. On substituting Eq. (5) into Eq. (9), the diagonal part of the density matrix of the resonator after $N$ rounds of equally spaced evolution and measurement turns out to be
$\mathcal{D}[\rho_{a}(N\tau)]=\sum_{k=0}^{\infty}\dfrac{|\alpha_{k}|^{2}|\lambda_{k}^{(e)}(\tau)|^{2N}|k\rangle\langle k|}{P_{e}(N)}.$ (13)
The fidelity of the target Fock state $|n\rangle$ is found to be
$\mathcal{F}(N)=\langle n|\rho_{a}(N\tau)|n\rangle=\frac{|\alpha_{n}|^{2}|\lambda_{n}^{(e)}(\tau)|^{2N}}{\sum_{k}|\alpha_{k}|^{2}|\lambda_{k}^{(e)}(\tau)|^{2N}}.$ (14)
To protect the population on the state $|n\rangle$, the relevant population-reduction factor has to satisfy $|\lambda^{(e)}_{n}(\tau)|^{2}=1$. Due to Eq. (5), the evolution period of a single round $\tau$ can be set as
$\tau=l\tau_{n}^{(e)}=l\frac{\pi}{\Omega^{(e)}_{n}},$ (15)
where $l\subseteq\mathbb{N}^{+}$. If the irrelevant population-reduction factors $|\lambda^{(e)}_{k}(l\tau_{k}^{(e)})|^{2}$ for all the significantly occupied Fock states $|k\rangle$, $k\neq n$, are always less than unit, then the resonator state will gradually approach the desired $|n\rangle$ as the measurement number $N$ increases. The filter effect can be further amplified by the success probability $P_{e}(N)<1$ in the denominator of Eq. (13). Therefore the density matrix of the resonator mode becomes
@@
in the ideal large-$N$ limit. Also, it indicates that the success probability converges to the initial population of the target Fock state:
@@
Figure 2: Population-reduction factors for generating the Fock state $|5\rangle$ as a function of the Fock index $k$ after a single evolution-measurement cycle with various cycle periods $\tau$ determined by different detuning $\delta$ and integer $l$. Here $g=0.05\omega_{e}$.
In Fig. 2 we present the population-reduction factors for generating the Fock state $|n=5\rangle$ as a function of the Fock index $k$ after one evolution-measurement cycle with optimized $\tau$. It is found with a fixed $l$ that a nonzero detuning $\delta$ gives rise to a smaller reduction rate for the neighboring states around $|5\rangle$ than a vanishing detuning. It will reduce our protocol’s efficiency. Besides the target state $|n\rangle$, also the populations on the Fock states $|n_{j}^{(l)}\rangle$ remain invariant when
$n_{j}^{(l)}=\frac{j^{2}}{l^{2}}(n+1)+\left(\frac{j^{2}}{l^{2}}-1\right)\frac{\delta^{2}}{4g^{2}}-1$ (18)
with $j\subseteq\mathbb{N}^{+}$ due to Eqs. (5) and (15). Under the same $\delta$, the solution set with $l\neq 1$ covers that with $l=1$ as shown by the red dashed line and green dotted line in Fig. 2. To perform strategies for Fock-state generation with varying $\tau$ or $l$ in the sequence of measurements, we then focus on the intersection of the solution set with $l=1$ and that with $l\neq 1$. Thus, in the following, $n_{j}^{(l)}$ can be briefly labeled as $n_{j}$ and it is readily found that $n=n_{1}$. More generally, a sufficiently large number of evolution-measurement cycles with $\tau$ in Eq. (15) gives rise to a stabilized subspace $\mathcal{V}_{n}^{(e)}$ spanned by $\{\left|n_{j}\right\rangle|j\subseteq\mathbb{N}^{+}\}$, where both the populations and coherence are under protection. To enhance the final success probability of the target Fock state, the initial coherent state $\left|\alpha\right\rangle$ of the resonator can be prepared with $|\alpha|^{2}=n$ such that $|n\rangle$ is the maximally occupied Fock state and the populations on $|n_{j\geq 2}\rangle$ are negligible. For example, when $n=n_{1}=5$ and $\delta=0$, the population on the next unwanted yet protected Fock-state $n_{2}=23$ is found to be $p_{n_{2}}\sim 10^{-5}$, about four orders lower than $p_{n_{1}}$.
Figure 3: Population-reduction factors as a function of the Fock index $k$ after $N$ evolution-measurement cycles with various evolution periods $\tau$ indicated by $l$. The target states are (a) $|n=5\rangle$ and (b) $|n=10\rangle$. The black solid lines indicate a single round with $\tau=\tau_{n}^{(e)}$ ($l=1$). The blue dotted lines, red dashed lines, and yellow dash-dotted lines indicate $N=5$ rounds with $\tau=\tau_{n}^{(e)}$, $\tau=2\tau_{n}^{(e)}$ ($l=2$), and $\tau=3\tau_{n}^{(e)}$ ($l=3$), respectively. Here $g=0.05\omega_{e}$ and $\delta=0$.
@@
The efficiency for generating the target state is determined by the profile of the population-reduction factors $|\lambda^{(e)}_{k}(l\tau_{n})|^{2N}$ in the Fock space as shown in Figs. 3(a) and 3(b), where $n=5$ and $10$, respectively. These factors are evaluated for various periods $\tau$ of free-evolution (indicated by $l$) after $N$ rounds of evolution and measurement. The black solid lines in Fig. 3 demonstrate the effect of a single projective measurement, confirming that the population-reduction factors of the irrelevant Fock states $k\neq n_{j\geq 1}$ satisfy $|\lambda_{k}^{(e)}(\tau)|^{2}<1$. Further, the comparison between the blue dotted lines and the black solid lines indicates that the populations in the neighboring regions of $|n_{j}\rangle$ can be significantly suppressed by a sufficient number of measurements.
@@
Using multiples of $\tau$ (see the red dashed lines for $l=2$ and the yellow dashed-dot line for $l=3$), it is interesting to find that, on the one hand, they can intensify the population-reduction effect surrounding the target state $|n_{1}\rangle$ with a highly concentrated population-protection region; yet on the other hand, more protected (although not strictly) Fock states arise, e.g., see the high profiles between $|n_{1}=5\rangle$ and $|n_{2}=23\rangle$ in Fig. 3(a). Those results suggest a compromise or hybrid strategy labeled with $\mathcal{S}_{l}^{(q)}$ that filters out the populations distributed on the states other than $|n_{j}\rangle$ by using $\tau=\tau^{(e)}_{n}$ in the first $q$ rounds of evolution and measurement and then promotes the state-generation efficiency by using $\tau=l\tau^{(e)}_{n}$ with $l>1$ in the rest rounds. Often we take $l=2$ or $l=3$ since the advantage from an even larger $l$ is not significant. Accordingly, the uniform strategy with an invariant period $\tau=\tau^{(e)}_{n}$ can be labeled with $\mathcal{S}_{1}^{(\infty)}$.
@@
Figure 4: Comparison of various strategies in Fock-state generation of (a) and (c) $|n=5\rangle$ and of (b) and (d) $|n=10\rangle$ as a function of the measurement number $N$. (a) and (b) Target-state fidelity. (c) and (d) Success probability, where the gray dotted lines indicate the initial population of $|n\rangle$. The system parameters are the same as in Fig. 3.
@@
In Fig. 4 we compare the efficiencies for generating the Fock state between the uniform strategy and the hybrid strategy $\mathcal{S}^{(5)}_{l}$ with $l=2,3$ in terms of the target-state fidelity $\mathcal{F}(N)$ and the success probability $P_{e}(N)$, whose definitions can be found in Eqs. (14) and (11), respectively. Within the current hybrid strategy, the first $q=5$ rounds of the evolution and measurement use $\tau=\tau_{n}^{(e)}$ and the remaining rounds use $\tau=l\tau_{n}^{(e)}$. It is found that the hybrid strategy overwhelms the uniform one in the state fidelity within a limited number of measurements. In addition, $l=3$ gives rise to the best performance. In Fig. 4(a) the fidelity $\mathcal{F}$ of the target Fock state $|5\rangle$ can reach $99.6\%$ within $15$ measurements when $l=3$. In contrast, $\mathcal{F}$ reaches only $68.6\%$ within $20$ measurements by the uniform strategy. In Fig. 4(b) it is found that a higher Fock state $|10\rangle$ could be obtained by merely $30$ measurements in the strategy $\mathcal{S}^{(5)}_{3}$ with a fidelity over $98.4\%$. A different strategy yields the same asymptotic value of the success probability, i.e., the population of $|n\rangle$ in the initial coherent state, which can be obtained by Eq. (17) with $|\alpha|^{2}=n$. In both Figs. 4(c) and 4(d), a larger $l$ gives rise to a faster convergence.
@@
Figure 5: Dependence of (a) fidelity and (b) success probability on the measurement number $N$ under the strategy $\mathcal{S}_{3}^{(5)}$ for various target state. The system parameters are the same as in Fig. 3.
@@
In Fig. 5 we generate different Fock states under the same hybrid strategy $\mathcal{S}_{3}^{(5)}$. The initial coherent states are accordingly set by $|\alpha|^{2}=n$. It is found in Fig. 5(a) that more rounds of evolution-measurement cycles are required to prepare a higher Fock states. On the one hand, a larger $n$ indicates a broader population distribution in Fock space and a smaller overlap between the corresponding coherent state and the target state; and on the other hand, a larger $n$ leads to a smaller difference in Rabi frequencies associated with $|n\rangle$ and $|n\pm 1\rangle$ and then more measurements are required to discriminate them. For example, it requires only $6$ rounds of evolution and measurement to enhance the fidelity of the target Fock state $|2\rangle$ to $99\%$, yet about $23$ rounds to prepare $|8\rangle$ with almost the same fidelity. Figure 5(b) shows that the asymptotic behavior of the success probability is insensitive to $n$. It declines rapidly for the first few rounds and approaches the steady value after about $N=10$ rounds. During the state-generation process, both target resonator and ancillary qubit cannot be completely isolated from the surrounding environment. The attainable fidelities of the target Fock states will be limited due to the coupling between the composite system and the environment (assumed to be at zero temperature). Under the Born-Markov approximation, the evolution of the composite system can be described by the Lindblad master equation:
$\dot{\rho}(t)=$ $-i[H^{\prime},\rho(t)]+\gamma\mathcal{L}[|g\rangle\langle e|]\rho(t)$ (19)
$+\kappa\mathcal{L}[a]\rho(t)+\gamma_{\phi}\mathcal{L}[|e\rangle\langle e|]\rho(t),$
where $\gamma$ and $\kappa$ represent the decay rates of the ancillary qubit and the target resonator, respectively, and $\gamma_{\phi}$ is the dephasing rate of the qubit. For simplicity, we set $\gamma=\gamma_{\phi}=2\kappa$. The Lindblad superoperator $\mathcal{L}[\mathcal{O}]$ is defined as
$\mathcal{L}[\mathcal{O}]\rho=\frac{1}{2}(2\mathcal{O}\rho\mathcal{O}^{\dagger}-\mathcal{O}^{\dagger}\mathcal{O}\rho-\rho\mathcal{O}^{\dagger}\mathcal{O}),$ (20)
Figure 6: Performance of our protocol in generating $|n=5\rangle$ under the strategy $\mathcal{S}_{3}^{(5)}$ in the presence of decoherence with various cooperativity $g^{2}/\kappa\gamma$. (a) Fidelity of the target state. (b) Success probability. (c) Population of the nearest-neighboring state $|4\rangle$. The system parameters are the same as in Fig. 3.
In Fig. 6(a), we present the fidelity for varying system cooperativity $g^{2}/\kappa\gamma$ for generating the target Fock state $|5\rangle$. In the presence of decoherence measured by the cooperativity $g^{2}/\kappa\gamma=10^{6}$, $10^{5}$, and $10^{4}$, the final fidelity of the target Fock state $|5\rangle$ can be stabilized at about $97\%$, $92\%$, and $76\%$, respectively, after a sufficient number of rounds of evaluation and measurement. It is the trade-off between the projective measurements and the decay from the desired Fock state to the lower states. In contrast to the fidelity, the success probability of measuring the ancillary qubit in its excited state $|e\rangle$ gradually decreases with $N$ due to the decoherence, as shown in Fig. 6(b). To show the underlying reasons, we present the population dynamics of the Fock state $|4\rangle$ as the nearest-neighboring lower state of the target state in Fig. 6(c). It is found that through the decay channel, $P_{4}$ increases in the first few rounds and then decreases with more measurements. However, the decoherence effect reduces the capability of our protocol to discriminate $|n-1\rangle$ from the target state $|n\rangle$, which results in a finite $P_{4}$ especially for a low system cooperativity.
@@ III.2 Superposed Fock-state generation
Beyond the existing works for Fock state generation, our protocol can be used to prepare a superposed Fock state such as
@@
with the desired $n$, $c_{0}$, and $c_{n}$, rather than a single Fock state Harel and Kurizki (1996) or a mixture of Fock states Delakouras et al. (2023). This target can be achieved by preparing the ancillary qubit in its ground state $|g\rangle$ and performing a sequence of projective measurements $M_{g}=|g\rangle\langle g|$ on the ancillary qubit. Due to Eqs. (4) and (5), the effective time-evolution operator for the resonator under the resonant condition $\delta=0$ and an invariant $\tau$ is now written as,
@@
In this case, the population in the vacuum state $\left|0\right\rangle$ of the resonator is always under protection due to the decoupling of the state $\left|g0\right\rangle$ from the other composite states in the time evolution. Using Eqs. (5) and (9), the diagonal part of the density matrix of the resonator after $N$ rounds of equally spaced evolution and measurement can be expressed as
@@
From Eqs. (21), (23), and (25), the fidelity $\mathcal{F}_{\pm}$ about the target state $|\psi_{\pm}\rangle$ after $N$ equally-spaced evolution-and-measurement rounds can be obtained as
$\mathcal{F}_{\pm}=$ $\bigg{[}|c_{0}|^{2}p_{0}+|c_{n}|^{2}p_{n}\cos^{2N}\left(\sqrt{n}g\tau\right)$
±$\pm 2{\rm Re}(C_{n0}c_{0}c_{n}^{*})\cos^{N}\left(\sqrt{n}g\tau\right)\bigg{]}/P_{g}(N).$ (26)
Figure 7: (a) Fidelities $\mathcal{F}_{\pm}$ for the target superposed state $(|0\rangle\pm|5\rangle)/\sqrt{2}$ or their sum $\mathcal{F}_{+}+\mathcal{F}_{-}$ as a function of $N$ under the uniform strategy $\mathcal{S}_{1}^{(\infty)}$ with $\tau=\tau_{5}^{(g)}$ and a hybrid strategy $\mathcal{S}_{3}^{(5)}$ that modifies the round period to be $\tau=3\tau_{5}^{(g)}$ after five rounds of evolution and measurement. (b) Success probability of finding the qubit in its ground state after $N$ measurements. The system parameters are the same as in Fig. 3.
Both populations in the ground state $|0\rangle$ and the Fock state $|n\rangle$ and their coherence should be under protection to generate the target superposed state $|\psi_{\pm}\rangle$ in Eq. (21). To this end, the evolution periods for each round of evolution and measurement are chosen as ($\delta=0$)
@@
where $l\in\mathbb{N}^{+}$. With $l=1$, a stabilized subspace under full protection emerges over multiple measurements. It reads
$\mathcal{V}^{(g)}_{n}={\rm span}\{|0\rangle,|n\rangle,|2^{2}n\rangle,|3^{2}n\rangle,...\}.$ (28)
The ratio of the probability amplitudes $c_{0}/c_{n}$ is determined by $\alpha_{0}/\alpha_{n}$. Then one can inversely find that the initial coherent state of the resonator is $|\alpha\rangle$ with $\alpha=(c_{n}\sqrt{n!}/c_{0})^{1/n}$. Similar to the Fock-state generation in Sec. III.1, the initial populations on the Fock-state elements in the stabilized states $\{|j^{2}n\rangle|j\geqslant 2\}$ are negligible. Consequently, the diagonal and off-diagonal parts of the density matrix of the resonator in the large-$N$ limit read
$\mathcal{D}[\rho_{a}(N\tau)]$ $\rightarrow e^{-|\alpha|^{2}}\left(|0\rangle\langle 0|+\frac{|c_{n}|^{2}}{|c_{0}|^{2}}|n\rangle\langle n|\right)\bigg{/}P_{g}(N),$ (29)
$\mathcal{C}[\rho_{a}(N\tau)]$ $\rightarrow e^{-|\alpha|^{2}}\left[\frac{c_{n}}{c_{0}}(-1)^{N}|n\rangle\langle 0|+{\rm H.c.}\right]\bigg{/}P_{g}(N),$ (30)
@@
$P_{g}(N)\rightarrow p_{0}+p_{n}=\exp\left[-\left(\frac{|c_{n}|}{|c_{0}|}\sqrt{n!}\right)^{2/n}\right]\bigg{/}|c_{0}|^{2}.$ (31)
@@
Figure 8: Density matrix of the resonator after various rounds of evolution and measurement under (a) and (c) the uniform strategy or (b) and (d) the hybrid strategy $\mathcal{S}_{3}^{(5)}$ for (a) and (b) $N=10$ and (c) and (d) $N=11$. The target state is $(|0\rangle\pm|5\rangle)/\sqrt{2}$ and the other parameters are the same as in Fig. 3.